The sum of two sisters ages is 60. One sister is 8 years older than the other. How old are they? And what is the equation I could use to solve the problem?
Answer by ankor@dixie-net.com(22740)

The sum of two sisters ages is 60.
One sister is 8 years older than the other.
How old are they?
:
Use two equations
Two sisters a & b
:
write an equation for each statement:
:
"The sum of two sisters ages is 60."
a + b = 60
" One sister is 8 years older than the other."
a = b + 8
or
a - b = 8
:
We can use elimination now
a + b = 60
a - b = 8
--------------adding eliminates b
2a = 68
a = 68/2
a = 34 is the age of one sister
:
You can find b I am sure